# Kai Shu, Lu Cheng, Huan Liu

### v1

@@ 4.1.1 Data Source
While manually-labeled feedback is high-quality and accurately reflects human judgment preferences, it has limitations in terms of quantity and coverage. As a result, some researchers have turned to synthetic feedback as a data source for tuning judge LLMs. One approach in this direction relies on the judge LLMs themselves to generate the synthetic feedback. For example, Wu et al. (2024) construct pairwise feedback for judgment enhancement by prompting the policy LLMs to evaluate their own judgments.  Wang et al. (2024f) prompt the LLM to generate a “noisy” version of the original instruction and use the corresponding response to this corrupted instruction as the inferior response.  Wang et al. (2024a) prompt GPT-4-Turbo to generate multiple pieces of evidence based on the original evidence for each instance, categorizing them into completely irrelevant evidence, partially irrelevant evidence and highly related evidence to train a hallucination judgment LLMs.  Park et al. (2024) build OFFSETBIAS, a pairwise preference dataset that leverages GPT-4 to generate bad, off-topic and erroneous responses and perform difficulty filtering. For safety judging, Xie et al. (2024a) adopt GPT-4 as the classifier to map each data point to a predefined safety category to train an automated evaluator. Different from previous works, Li et al. (2024e) adopt GPT-4 to synthesize both pairwise and pointwise data to train a generative judge LLM. For pointwise data, they adopt a “divide-and-conquer” strategy, where two critiques are collected from GPT-4 for a single response, combined into a more comprehensive critique, and a final rating is provided Following them, Kim et al. (2024) use GPT-4 to augment the preference learning dataset with detailed human evaluation criteria and verbal feedback. In the multi-modal domain, Xiong et al. (2024b) propose LLaVA-Critic and adopt GPT-4o to generate reasons behind given scores or preference judgments for training data construction. Moreover, By harnessing both explicit human instruction and the implicit knowledge of GPT-4, Xu et al. (2023) fine-tune a judge LLM based on LLaMA, producing both a score for generated text and a human-readable diagnostic report.  Zhu et al. (2023) introduce JudgeLM and propose a comprehensive, large-scale, high-quality dataset containing task seeds, LLMs-generated answers, and GPT-4-generated judgments for fine-tuning high-performance judges.
@@ 4.1.2 Tuning Techniques
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024f), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024j) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.
@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023); Kim et al. ; Wang et al. (2024e); Murugadoss et al. (2024) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024e) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently, Dong et al. (2024) investigate the reliability of LLM-as-a-personalized-judge, providing persona as a part of principles for LLMs to make personalized judgments.
@@ 4.2.3 Multi-agent Collaboration
Accessing results from a single LLM judge may not be reliable due to various biases inherent in LLMs Wang et al. (2023c); Liusie et al. (2024); Zhu et al. (2023); Koo et al. (2023a); Liu et al. (2023c). To address this limitation, Li et al. (2023) introduced the Peer Rank (PR) algorithm, which considers each peer LLM’s pairwise preferences for all answer pairs and produces a final ranking of models. Building on this foundation, several architectures and techniques for multi-agent LLMs have emerged, including mixture-of-agent Zhang et al. (2023b), role play Wu et al. (2023), debating Chan et al. (2023), and voting Zhu et al. (2024b). Jung et al. (2024) proposed Cascaded Selective Evaluation, where less expensive models serve as initial judges, escalating to stronger models only when necessary Beigi et al. (2024a). Additionally, some works apply multi-agent collaboration for alignment data synthesis, leveraging multiple LLM judges to refine responses Arif et al. (2024) or provide more accurate pairwise feedback Li et al. (2024g).
@@ 4.2.5 Multi-turn Interaction
In evaluation, a single response may not provide enough information for an LLM judge to thoroughly and fairly assess each candidate’s performance. To address this limitation, multi-turn interactions are commonly adopted to offer a more comprehensive evaluation. Typically, the process begins with an initial query or topic, followed by dynamically interacting between the judge LLM and candidate models. Bai et al. (2023b) propose a multi-round setting where the evaluator takes on an interviewer role, posing increasingly sophisticated follow-up questions based on the models’ previous answers. Similarly, Yu et al. (2024) introduce KIEval, a Knowledge-grounded Interactive Evaluation framework, which novelly incorporates an LLM-powered interactor to enable dynamic, contamination-resilient assessments. Additionally, some approaches facilitate debates among candidates in a multi-round format. For example, Zhao et al. (2024c) design a framework where two LLMs engage in a multi-round peer battle around a query, allowing their true performance differences to surface. Moniri et al. (2024) propose an automated benchmarking system where LLMs debate, with the final assessment carried out by another LLM judge.
@@ 4.2.6 Comparison Acceleration
Among the various comparison formats in LLM-as-a-judge (e.g., point-wise and list-wise), pair-wise comparison is the most common approach for directly comparing two models or generating pair-wise feedback. However, when multiple candidates need to be ranked, this method can be quite time-consuming Zhai et al. (2024). To mitigate the computational overhead, Zhai et al. (2024) propose a ranked pairing method in which all candidates are first compared against a blank baseline response. Each candidate’s rank is then determined by how well they perform in comparison to the baseline. Zhu et al. (2024a) proposed a CoT like prompting technique to mitigate the positional bias by forcing the model to provide all pairwise ranking first, then summarize these pairwise rankings with a list. In addition, Lee et al. (2024) utilize a tournament-based approach Liu et al. (2023a); Zhao et al. (2023b) for rejection sampling during inference to speed up the pair-wise comparison. They construct a tournament tree where the leaf nodes represent sampled responses, and non-leaf nodes are selected based on the outcome of judgments between child nodes.
@@ 5.1.1 Open-ended Generation Tasks
Open-ended generation refers to tasks where the generated content is expected to be safe, accurate, and contextually relevant, though there isn’t a single “correct” answer. Such tasks include dialog response generation, summarization, story generation, and creative writing Badshah and Sajjad (2024); Bai et al. (2023a); Kumar et al. (2024). Unlike conventional metrics-based evaluation methods, LLM-as-a-judge provides a more nuanced, adaptable, and customized evaluation. As noted by Zheng et al. (2023), LLMs like GPT-4 perform comparably to humans when judging open-ended text generation. In practice, LLM-as-a-judge has been applied to evaluate outputs from a single model to compare outputs from multiple models in a competitive setting. For example, Gao et al. (2023) use ChatGPT to perform human-like summarization evaluation. Similarly, Wu et al. (2023) propose a comparison-based framework to let LLMs act as judges with multiple role-playing to evaluate the summarization quality in a specific dimension and generate its evaluations.
Modern LLMs are good at generating detailed and long-form responses. However, as the output length increases, so does the likelihood of hallucinations. To better understand this phenomenon, Cheng et al. (2023) and Zhang et al. (2024d) introduce an evaluation method that uses GPT-4 to judge whether generated outputs include logically structured yet nonsensical statements. Wang et al. (2024a) propose a critique-based judging system that evaluates hallucinations by selecting pertinent evidence and providing in-depth critiques. Beyond hallucinations, a significant concern is the generation of harmful (e.g., encouragement of self-suicide) and unsafe (e.g., guidance of illegal activity) responses by LLMs. Addressing this, Li et al. (2024f) introduce MD-Judge and MCQ-Judge for evaluating safety-related QA pairs, particularly focusing on queries crafted to provoke unsafe responses. This approach supports a seamless and reliable evaluation. However, an overly cautious stance toward unsafe queries can lead to excessive refusal responses, hindering normal functionality and negatively affecting user experience. To explore this issue, Xie et al. (2024a) conduct a meta-evaluation across various LLM-as-a-judge frameworks, assessing the refusal tendencies of current LLMs in response to potentially unsafe queries.
@@ 5.1.3 Emerging Tasks
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024j), debating judgment Liang et al. (2024a) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024g) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024d) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024c) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration.
Several works have also employed LLM-as-a-judge in specific domains or for particular attributes. Zhang et al. (2024f) propose a self-evaluation mechanism, judging responses’ factuality by generating question-answer pairs. They then utilize these self-annotated responses to fine-tune the model via the DPO algorithm for better factuality. In robotics, Zeng et al. (2024) iteratively update the reward function with the self-ranking responses from LLMs, boosting learning efficiency without human supervision. In the multimodal domain, Ahn et al. (2024) propose iterative self-retrospective judgment (i-SRT), which employs self-reflection to improve both response generation and preference modeling.

### v2

@@ 4.1.1 Data Source
While manually-labeled feedback is high-quality and accurately reflects human judgment preferences, it has limitations in terms of quantity and coverage. As a result, some researchers have turned to synthetic feedback as a data source for tuning judge LLMs. One approach in this direction relies on the judge LLMs themselves to generate the synthetic feedback. For example, Wu et al. (2024a) construct pairwise feedback for judgment enhancement by prompting the policy LLMs to evaluate their own judgments.  Wang et al. (2024f) prompt the LLM to generate a “noisy” version of the original instruction and use the corresponding response to this corrupted instruction as the inferior response.  Wang et al. (2024a) prompt GPT-4-Turbo to generate multiple pieces of evidence based on the original evidence for each instance, categorizing them into completely irrelevant evidence, partially irrelevant evidence and highly related evidence to train a hallucination judgment LLMs.  Park et al. (2024) build OFFSETBIAS, a pairwise preference dataset that leverages GPT-4 to generate bad, off-topic and erroneous responses and perform difficulty filtering. For safety judging, Xie et al. (2024a) adopt GPT-4 as the classifier to map each data point to a predefined safety category to train an automated evaluator. Different from previous works, Li et al. (2024e) adopt GPT-4 to synthesize both pairwise and pointwise data to train a generative judge LLM. For pointwise data, they adopt a “divide-and-conquer” strategy, where two critiques are collected from GPT-4 for a single response, combined into a more comprehensive critique, and a final rating is provided Following them, Kim et al. (2024) use GPT-4 to augment the preference learning dataset with detailed human evaluation criteria and verbal feedback. In the multi-modal domain, Xiong et al. (2024b) propose LLaVA-Critic and adopt GPT-4o to generate reasons behind given scores or preference judgments for training data construction. Moreover, By harnessing both explicit human instruction and the implicit knowledge of GPT-4, Xu et al. (2023) fine-tune a judge LLM based on LLaMA, producing both a score for generated text and a human-readable diagnostic report.  Zhu et al. (2023) introduce JudgeLM and propose a comprehensive, large-scale, high-quality dataset containing task seeds, LLMs-generated answers, and GPT-4-generated judgments for fine-tuning high-performance judges.
@@ 4.1.2 Tuning Techniques
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024f), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024n) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.
@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024b); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023); Kim et al. ; Wang et al. (2024e); Murugadoss et al. (2024); Li et al. (2024k, h) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024f) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently, Dong et al. (2024) investigate the reliability of LLM-as-a-personalized-judge, providing persona as a part of principles for LLMs to make personalized judgments.
@@ 4.2.3 Multi-agent Collaboration
Accessing results from a single LLM judge may not be reliable due to various biases inherent in LLMs Wang et al. (2023c); Liusie et al. (2024); Zhu et al. (2023); Koo et al. (2023a); Liu et al. (2023c). To address this limitation, Li et al. (2023) introduced the Peer Rank (PR) algorithm, which considers each peer LLM’s pairwise preferences for all answer pairs and produces a final ranking of models. Building on this foundation, several architectures and techniques for multi-agent LLMs have emerged, including mixture-of-agent Zhang et al. (2023b), role play Wu et al. (2023), debating Chan et al. (2023); Zhang et al. (2024e), and voting Zhu et al. (2024b). Jung et al. (2024) proposed Cascaded Selective Evaluation, where less expensive models serve as initial judges, escalating to stronger models only when necessary Beigi et al. (2024a). Additionally, some works apply multi-agent collaboration for alignment data synthesis, leveraging multiple LLM judges to refine responses Arif et al. (2024) or provide more accurate pairwise feedback Li et al. (2024i). Recnetly, Li et al. (2024l) propose the MATEval, where all agents are played by LLMs like GPT-4. The MATEval framework emulates human collaborative discussion methods, integrating multiple agents’ interactions to evaluate open-ended text.
@@ 4.2.5 Multi-turn Interaction
In evaluation, a single response may not provide enough information for an LLM judge to thoroughly and fairly assess each candidate’s performance. To address this limitation, multi-turn interactions are commonly adopted to offer a more comprehensive evaluation. Typically, the process begins with an initial query or topic, followed by dynamically interacting between the judge LLM and candidate models. Bai et al. (2023b) propose a multi-round setting where the evaluator takes on an interviewer role, posing increasingly sophisticated follow-up questions based on the models’ previous answers. Similarly, Yu et al. (2024b) introduce KIEval, a Knowledge-grounded Interactive Evaluation framework, which novelly incorporates an LLM-powered interactor to enable dynamic, contamination-resilient assessments. Additionally, some approaches facilitate debates among candidates in a multi-round format. For example, Zhao et al. (2024c) design a framework where two LLMs engage in a multi-round peer battle around a query, allowing their true performance differences to surface. Moniri et al. (2024) propose an automated benchmarking system where LLMs debate, with the final assessment carried out by another LLM judge.
@@ 4.2.6 Comparison Acceleration
Among the various comparison formats in LLM-as-a-judge (e.g., point-wise and list-wise), pair-wise comparison is the most common approach for directly comparing two models or generating pair-wise feedback. However, when multiple candidates need to be ranked, this method can be quite time-consuming Zhai et al. (2024). To mitigate the computational overhead, Zhai et al. (2024) propose a ranked pairing method in which all candidates are first compared against a blank baseline response. Each candidate’s rank is then determined by how well they perform in comparison to the baseline. Zhu et al. (2024a) proposed a CoT like prompting technique to mitigate the positional bias by forcing the model to provide all pairwise ranking first, then summarize these pairwise rankings with a list. In addition, Lee et al. (2024) utilize a tournament-based approach Liu et al. (2023a); Zhao et al. (2023c) for rejection sampling during inference to speed up the pair-wise comparison. They construct a tournament tree where the leaf nodes represent sampled responses, and non-leaf nodes are selected based on the outcome of judgments between child nodes.
@@ 5.1.1 Open-ended Generation Tasks
Open-ended generation refers to tasks where the generated content is expected to be safe, accurate, and contextually relevant, though there isn’t a single “correct” answer. Such tasks include dialog response generation, summarization, story generation, and creative writing Badshah and Sajjad (2024); Bai et al. (2023a); Kumar et al. (2024); Zeng et al. ; Yang et al. (2024). Unlike conventional metrics-based evaluation methods, LLM-as-a-judge provides a more nuanced, adaptable, and customized evaluation. As noted by Zheng et al. (2023), LLMs like GPT-4 perform comparably to humans when judging open-ended text generation. In practice, LLM-as-a-judge has been applied to evaluate outputs from a single model to compare outputs from multiple models in a competitive setting. For example, Gao et al. (2023) use ChatGPT to perform human-like summarization evaluation. Similarly, Wu et al. (2023) propose a comparison-based framework to let LLMs act as judges with multiple role-playing to evaluate the summarization quality in a specific dimension and generate its evaluations.
Modern LLMs are good at generating detailed and long-form responses. However, as the output length increases, so does the likelihood of hallucinations. To better understand this phenomenon, Cheng et al. (2023) and Zhang et al. (2024d) introduce an evaluation method that uses GPT-4 to judge whether generated outputs include logically structured yet nonsensical statements. Wang et al. (2024a) propose a critique-based judging system that evaluates hallucinations by selecting pertinent evidence and providing in-depth critiques. Beyond hallucinations, a significant concern is the generation of harmful (e.g., encouragement of self-suicide) and unsafe (e.g., guidance of illegal activity) responses by LLMs. Addressing this, Li et al. (2024g) introduce MD-Judge and MCQ-Judge for evaluating safety-related QA pairs, particularly focusing on queries crafted to provoke unsafe responses. This approach supports a seamless and reliable evaluation. However, an overly cautious stance toward unsafe queries can lead to excessive refusal responses, hindering normal functionality and negatively affecting user experience. To explore this issue, Xie et al. (2024a) conduct a meta-evaluation across various LLM-as-a-judge frameworks, assessing the refusal tendencies of current LLMs in response to potentially unsafe queries. Besides, Yu et al. (2024a) introduce an LLM-based answer extractor to accurately pinpoint critical part of answer from the open-ended generation.
@@ 5.1.3 Emerging Tasks
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024j), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024d) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024c) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.
Several works have also employed LLM-as-a-judge in specific domains or for particular attributes. Zhang et al. (2024g) propose a self-evaluation mechanism, judging responses’ factuality by generating question-answer pairs. They then utilize these self-annotated responses to fine-tune the model via the DPO algorithm for better factuality. In robotics, Zeng et al. (2024) iteratively update the reward function with the self-ranking responses from LLMs, boosting learning efficiency without human supervision. In the multimodal domain, Ahn et al. (2024) propose iterative self-retrospective judgment (i-SRT), which employs self-reflection to improve both response generation and preference modeling.